### v2

# Competing LLM Agents in a Non-Cooperative Game of Opinion Polarisation

Abstract
We introduce a novel non-cooperative game to analyse opinion formation and resistance, incorporating principles from social psychology such as confirmation bias, resource constraints, and influence penalties. Our simulation features Large Language Model (LLM) agents competing to influence a population, with penalties imposed for generating messages that propagate or counter misinformation. This framework integrates resource optimisation into the agents’ decision-making process. Our findings demonstrate that while higher confirmation bias strengthens opinion alignment within groups, it also exacerbates overall polarisation. Conversely, lower confirmation bias leads to fragmented opinions and limited shifts in individual beliefs. Investing heavily in a high-resource debunking strategy can initially align the population with the debunking agent, but risks rapid resource depletion and diminished long-term influence.
Competing LLM Agents in a Non-Cooperative Game of Opinion Polarisation
Amin Qasmi2, Usman Naseem3, Mehwish Nasim1 1The University of Western Australia, 2Lahore University of Management Sciences, 3Macquarie University, Correspondence: mehwish.nasim@uwa.edu.au
1 Introduction and Background
The study of opinion dynamics, originating from efforts to understand how individuals modify their views under social influence (Kelman, 1958, 1961), has broad applications in areas such as public health campaigns, conflict resolution, and combating misinformation. Within social networks, opinions spread and evolve, influenced by various factors including peer interactions Kandel (1986), media exposure Zucker (1978), and group dynamics Friedkin and Johnsen (2011). Developing accurate models of these processes is essential not only for predicting trends like opinion polarisation (Tan et al., 2024) or consensus formation but also for crafting targeted interventions to mitigate harmful effects, such as the spread of misinformation or societal fragmentation (Hegselmann and Krause, 2015). Agent-based models (ABMs), simulating interactions among individual agents as proxies for humans, serve as valuable tools for examining the emergent properties of opinion dynamics. These models offer robust frameworks for analysing complex scenarios Deffuant et al. (2002); Mathias et al. (2016), evaluating strategies to reduce negative consequences, and potentially fostering constructive social influence by integrating explicit cognitive mechanisms into opinion-updating processes.
This work investigates how Large Language Models (LLMs) can model human-like opinion dynamics and influence propagation within social networks. Traditional ABMs often employ simplified rules that fail to capture the complexity of human communicative strategies. To address this limitation, we introduce a novel non-cooperative game framework where adversarial LLMs, one spreading misinformation and the other countering it, interact. This work introduces a non-cooperative game where LLM agents engage in adversarial interactions to model misinformation spread and countering. Unlike prior studies Wang et al. (2025); Chuang et al. (2024) on passive opinion evolution and nudging, it focuses on resource-constrained influence operations and debunking effectiveness in competitive environments.
We pose the following research questions:
  1. RQ1
    What are the emergent behaviors in networks of agents influenced by competing LLMs?
  2. RQ2
    How does the competition between LLM agents shape the evolution of opinion clusters over time, also known as echo-chambers?
Figure 1: In each round, the current team (Red/Blue) generates a message, which is then assigned a potency value by the judge. The network is updated as per the BCM Model’s update equation. In the next round, the other team receives the results of their opponent’s message in the previous round, as well as the results of their last message.
2 Methodology
We use LLMs to simulate the propagation and debunking of misinformation on social media within a non-cooperative game framework.
Scenario: Our scenario is strategically designed to reflect the asymmetric nature of contested information environments, specifically highlighting the challenges faced by the "Blue team" (i.e., those countering misinformation). This framework mirrors adversarial dynamics commonly modelled in serious games or wargames, particularly in the context of cybersecurity. The "Red Team" and "Blue Team" construct Paul, Christopher and Connable, Ben and Welch Jonathan and Rosenblatt, Nate and McNeive, Jim (2021), familiar in cybersecurity practices (as detailed in NIST’s Glossary National Institute of Standards and Technology (2015) (NIST)), is adapted to our simulation. The system comprises two LLM-based agents with opposing objectives: the Red Agent (adversarial agent) disseminates misinformation, while the Blue Agent (debunking agent) counters it and aims to restore trust. These agents operate within a directed network of neutral agents, termed Green Nodes, representing individuals in a population. Figure 1 shows the structure of our non-cooperative game.
Agent Roles and Mechanics: The simulation incorporates the following agent roles:
Red Agent: The Red Agent aims to amplify doubt and confusion by generating misinformation messages of varying potency. Messages of higher potency incur penalties through rejection, reflecting real-world scenarios where informed populations are sceptical of, and less susceptible to, high-strength misinformation.
Blue Agent: The Blue Agent counters the misinformation spread by the Red Agent, but operates under a resource constraint. High-potency counter-messages incur a greater resource cost. Therefore, the Blue Agent must strategically manage its available resources.
Judge Agent: To prevent agents from assigning arbitrarily high potencies to their messages, a Judge Agent assigns potencies based on specified criteria:
  • Clarity:Is the message clear and well-articulated?
  • Evidence: Does the message provide credible evidence or logical reasoning?
  • Relevance: Does the message effectively address the misinformation?
  • Impact: Does the message effectively persuade or influence the target audience?
Simulation settings: The simulation begins with $n$ nodes, of which $x$ are initially aligned with the Red Agent’s misinformation (pro-conspiracy), $y<n-x$ are aligned against it (anti-conspiracy), and the remaining $z=n-x-y$ are neutral. Both Red and Blue Agents generate messages, the potencies of which are determined by the Judge Agent.
Opinion Modeling: We use the Bounded Confidence Model (BCM) Mathias et al. (2016) to simulate opinion dynamics. In the BCM, a node updates its opinion if the difference between its opinion and that of a neighbouring node is less than a threshold (the confirmation bias value, $\mu$). The opinion update condition and formula are summarised below:
for all $n\in N$ do
     for all $m\in\text{Neighbours}(n)$ do
         if $|O_{m}-O_{n}|<\mu$ then
              $O_{m}\leftarrow O_{m}+\mu(O_{m}-O_{n})$
         end if
     end for
end for
where $N$ is the set of all nodes, $O_{n}$ is the opinion of node $n$, $O_{m}$ is the opinion of neighbour $m$, and $\text{Neighbours}(n)$ returns the neighbours of $n$. Opinions range between [-1, 1]. Nodes with opinions less than -0.5 are considered aligned with the Blue Agent, those greater than 0.5 with the Red Agent, and those in between are neutral.
Topics Classification: Topics for misinformation include serious debates and popular conspiracy theories (e.g., "The Earth is Flat") as well as more frivolous claims (e.g., "The Moon is made of cake").
Models: Our study employs GPT-4O and 4O-MINI as judges Hurst et al. (2024); OpenAI (2024). Experiment A compares Mixtral-8x7B-Instruct with Gemma-2-9b Jiang et al. (2024); Team et al. (2024b), while experiment B evaluates Mixtral-8x7B-Instruct against Gemini 1.5 Flash-8b Team et al. (2024a). Lastly, experiment C contrasts Gemma-2-9b with Gemini 1.5 Flash-8b.
Reinforcement learning: Reinforcement Learning from Human Feedback (RLHF) has proven effective in aligning LLMs with human preferences Kaufmann et al. (2023). In our simulations, agents receive feedback after each round, including their last message and metrics such as the percentage of followers gained or lost. This feedback allows them to refine their messaging strategies.
3 Simulations and Evaluation
We conducted simulations of 100 rounds for each experiment. The network configuration remained constant across all simulations: a directed network of 50 nodes, with 40% (20 nodes) initially aligned with the Blue Agent (anti-conspiracy) and 20% (10 nodes) aligned with the Red Agent (pro-conspiracy). This initial distribution reflects the observation that conspiracy theorists typically represent a minority within social media populations. The Blue Agent was initialised with a resource value of 100 and an influence factor (base potency) of 0.6, while the Red Agent’s influence factor was set to 0.5. We explored three Bounded Confidence Model (BCM) thresholds ($\mu$): 0.3, 0.7, and 0.9.
To investigate the impact of increased resource investment in debunking, we conducted a further set of experiments (using $\mu=0.9$) where the Blue Agent generated highly potent messages during the first 20 rounds (10 messages per agent). During these initial rounds, the Blue Agent’s messages had their base potency scaled by a factor of 1.2, with the final assigned potency capped at 1.0 (i.e., $\min(\text{potency}\times 1.2,1.0)$). This simulated a "high-resource" debunking strategy.
All simulations were run on a local machine having an Intel(R) Core(TM) i7-1355U 13th Gen CPU and 32 Gigabytes of RAM. While the machine was equipped with an integrated GPU (Intel Iris Xe Graphics (shared 15.8GB memory)), no dedicated GPU acceleration was used in this project. For all LLMs, the temperature was kept at 0.5, top_p at 1.0, and max_tokens were set to 100. The results of these simulations are presented in section 4.
Metrics: We employ the following metrics to analyse the results of our simulations.
Polarisation refers to the division of individuals or groups into opposing factions, leading to the reinforcement of extreme views within social networks. Network polarisation ($P$) is calculated Chitra and Musco (2020) as:
$P=\frac{1}{N}\sum_{n\in\mathcal{V}}(O_{n}-\bar{O})^{2}$ (1)
where:
  • $N$ is the total number of nodes in the network.
  • $\mathcal{V}$ is the set of all nodes.
  • $O_{n}$ is the opinion of node $n$.
  • $\bar{O}$ is the mean opinion across all nodes.
Judge’s Agreement: To assess the consistency of the potencies assigned by Judge Agent, we used two Judge Agents in each round to independently generate potency values for the same messages. All generated potency values were recorded, and inter-rater agreement was measured using two metrics: Intraclass Correlation Coefficient (ICC) and Krippendorff’s Alpha. ICC values range from 0 to 1, where values closer to 1 indicate strong agreement and values closer to 0 indicate poor agreement. Krippendorff’s Alpha ranges from -1 to 1, with higher values indicating stronger agreement.
BCM Threshold Experiment ICC Value Krippendorff’s Alpha
0.3 A 0.495 0.420
B 0.085 0.035
C 0.485 0.435
0.7 A 0.625 0.560
B 0.615 0.590
C 0.540 0.470
0.9 A 0.490 0.395
B 0.420 0.400
C 0.575 0.540
Table 1: Across Topic Average Intraclass Correlations (ICC) and Krippendorff’s Alpha for Different BCM Thresholds. The values represent the agreement between the agents while assigning potencies to messages.
Figure 2: (a) & (c) show population opinion percentages and average polarisation across topics for BCM thresholds 0.3, 0.7, and 0.9 over 100 rounds with models A(—), B(- - -), and $\dots$). Red and Blue colors indicate population’s alignment with adversarial and debunking agents respectively . Figures (b) & (d) present opinions and polarisation for BCM threshold 0.9 over 50 rounds for the same experiments.
4 Results and Discussion
RQ1 explores how adversarial interactions between competing LLM-driven agents (representing misinformation and counter-misinformation) influence collective opinion dynamics. Specifically, we examine the role of cognitive biases (represented by the BCM threshold) in shaping the stability and evolution of opinion alignment. Figure 2 summarises our findings.
Figure LABEL:fig:opinions illustrates the evolution of agent alignment over time for different BCM thresholds. At a low threshold ($\mu=0.3$), the opinion landscape becomes highly fragmented, with the Blue Agent’s alignment stagnating at its initial 40%. In contrast, at moderate and high thresholds ($\mu=0.7$ and $\mu=0.9$), the Blue Agent’s alignment increases to 45% and 50%, respectively. The Red Agent’s alignment also increases with higher thresholds, rising from 20% to 25%, 35%, and 38% at thresholds of 0.3, 0.7, and 0.9, respectively. This demonstrates that, without resource constraints, accumulating support is more feasible. Furthermore, the early rounds of interaction appear crucial in shaping long-term opinion trajectories, highlighting the strategic importance of early influence.
Figure LABEL:fig:polarisations shows the corresponding polarisation trends. A low threshold ($\mu=0.3$) results in only a marginal increase in polarisation ($\sim$40%), while higher thresholds ($\mu=0.7$ and $\mu=0.9$) lead to substantially higher polarisation levels ($\sim$70and $\sim$80%, respectively). These results align with the BCM update algorithm (Section 2) and the polarisation calculation (equa 1). With a small BCM threshold, agents only update their opinions if they are already closely aligned, resulting in multiple localised opinion clusters rather than a single consensus. Consequently, polarisation remains moderate as divergence occurs within sub-clusters. However, at high BCM thresholds, interactions occur more frequently across a broader range of opinions, amplifying extreme positions. As observed in Figure LABEL:fig:opinions (for $\mu=0.9$), nearly 90% of agents become strongly aligned with either the Red or Blue Agent, reflecting this sharp increase in polarisation.
RQ2 investigates optimal strategies for the Blue Agent to effectively counter misinformation under resource constraints. Specifically, we analyse the impact of an aggressive early-game approach, where high-potency debunking messages are deployed at a substantial resource cost. Due to the rapid resource depletion associated with this strategy, these simulations were limited to 50 rounds.
Figure LABEL:fig:opinions2 shows that this aggressive strategy enabled the Blue Agent to surpass the 50% alignment threshold, reaching a peak of 52%. Importantly, all three experimental conditions (A, B, and C) exhibited higher maximum alignment compared to the previous strategies, suggesting that an initial surge of high-potency messages leads to a greater overall shift towards the Blue Agent’s perspective.
Figure LABEL:fig:polarisations2 shows a sharp increase in polarisation during the first 20 rounds (corresponding to the high-resource debunking period), followed by a gradual convergence. This indicates that while an aggressive approach initially amplifies divisions, it eventually stabilises as the influence of misinformation diminishes. These findings highlight the trade-off between immediate impact and long-term sustainability in misinformation counter-strategies, emphasizing the importance of energy management in prolonged engagements.
5 Conclusions and Future Work
This study examines opinion polarization in a non-cooperative game with adversarial LLMs spreading and countering misinformation. Higher BCM thresholds enhance faction alignment but intensify societal polarization. We identify a trade-off between immediate impact and sustainability: high-impact interventions deplete resources quickly, while frequent interactions may deepen polarization. These findings inform LLM-driven influence operations and suggest future research on adaptive agents and real-world network integration.
Limitations
The study is based on simulated interactions rather than real-world datasets from social media or online discourse. Validating findings with empirical data would enhance their applicability. The study relies on the BCM, which, while effective, does not capture more complex psychological and social dynamics influencing opinion formation, such as emotional contagion, identity-based biases, or network homophily.
Ethical Statement
This study, involving the simulated generation of misinformation and counter-misinformation, necessitates careful ethical considerations. To prevent potential misuse, the specific prompts used to generate misinformation via LLMs cannot be disclosed. Disclosure could inadvertently facilitate real-world misinformation spread. Mitigation strategies included containing all generated content within a closed experimental environment, focusing the research objective on analysis and countermeasure development (not propagation), and ensuring that any released findings emphasise generalisable insights rather than specific prompt engineering techniques. We underscore that responsible misinformation modelling research is paramount, ensuring that the development of countermeasures does not contribute to the problem itself.